If the first differences of a sequence are a constant 3 and the second term is 8 find the first 5 terms of the sequence.
PLEASE SOLVE IF POSSIBLE.
AND COULD YOU PLEASE EXPLAIN HOW
TO SOLVE SO I UNDERSTAND.
THANK YOU SO MUCH!:]

Found 2 solutions by checkley71, stanbon:
Answer by checkley71(8403)

what you have is the sequence 5+3=8+3=11+3=14+3=17 so the sequence is 5,8,11,14,17

Answer by stanbon(48567)

If the first differences of a sequence are a constant 3 and the second term is 8 find the first 5 terms of the sequence.
------------
The information you were given shows:
1st term = ?
2nd term = 8
================
8 - ? = 3
? = 8-3
? = 5
------------------------------------
So, the sequence looks like:
5,8,11,14,17
==================
Cheers,
Stan H>